Find the eccentricity of earth`s orbit,if the minimum distance of the sun from the earth is 147.5 million kms and maximum distance is 152.5 million kms.
Answer by lwsshak3(11628)

Find the eccentricity of earth`s orbit,if the minimum distance of the sun from the earth is 147.5 million kms and maximum distance is 152.5 million kms.
**
Assuming the Earth's orbit around the sun is elliptical:
Consider the standard form of an equation for an ellipse: (x-h)^2/a^2+(y-k)^2/b^2=1, a>b, with (h,k) being the (x,y) coordinates of the center.
For given problem:
center: (0,0)
a=maximum distance of the sun from the earth= 152.5 million kms.
b=minimum distance of the sun from the earth= 147.5 million kms
c^2=a^2-b^2=(152.5)^2-(147.5)^2=1500
c=√1500≈38.73
eccentricity=c/a=38.73/152.5≈0.254